Given the following array index of a 5-based array implementation (indexes start at 5), representing an array that has the dimensions 20 x 20 (20 rows of 20 elements each), what is the linear (1D) address or index for the array element below?

Array1[6][7]

Answer is 22. Please show steps
Answer #1

index for the array element Array1[6][7] is
= 6*20 + 7
= 120 + 7
= 127

In 5-based array implementation
index for the array element Array1[6][7] is
= (6-5)*20 + (7-5)
= (1)*20 + (2)
= 20 + 2
= 22